OpenStudy (raffle_snaffle):

A sample of an ideal gas has its volume doubled while its temperature remains constant. If the original pressure was 100 torr, what is the new pressure?

OpenStudy (raffle_snaffle):

@wio

OpenStudy (raffle_snaffle):

@Abhisar

OpenStudy (raffle_snaffle):

@ganeshie8

OpenStudy (abhisar):

P\(\propto\)\(\huge\frac{1}{v}\)

OpenStudy (raffle_snaffle):

okay thanks.

OpenStudy (abhisar):

So u can use \(\huge\frac{P_1}{P_2}\) = \(\huge\frac{V_2}{V_1}\)

OpenStudy (raffle_snaffle):

i used that equation...

OpenStudy (raffle_snaffle):

solved for p_2

OpenStudy (raffle_snaffle):

my bad i used a different equation

OpenStudy (abhisar):

yes, P\(_1\) is given 100 torr, V\(_2\) is given 2*V\(_1\)

OpenStudy (abhisar):

Find P\(_2\)

OpenStudy (raffle_snaffle):

How did you know that you were suppose to use the equation you showed me

OpenStudy (abhisar):

From combined gas law i.e PV=nRT

OpenStudy (abhisar):

n, R & T are constant so P \(\propto\) \(\huge\frac{1}{v}\)

OpenStudy (abhisar):

if volume is doubled then pressure will become half

OpenStudy (raffle_snaffle):

a linear relationship? 

OpenStudy (abhisar):

yes

OpenStudy (raffle_snaffle):

okay thank you

OpenStudy (raffle_snaffle):

do you want to see my work?

OpenStudy (abhisar):

sure

OpenStudy (raffle_snaffle):